Your shadow measures 8 feet. Your father's shadow is 9 feet. If you are 5.5 feet, how tall is your father in inches? (round to nearest inch.)
Answer by Shana-D77(132)

Set up a proportion:
kid/dad = kid/dad
8/9 = 5.5/dad
Cross multiply to solve proportions:
8*dad = 9*5.5
8*dad = 49.5
dad = 6.1875 feet
They want inches so multiply by 12:
74.25 inches
They want "nearest inch" so:
Dad is 74 inches tall.